How much work is required to carry a 0.02μC charge from point r=40cm to that at r= 5cm


1
Expert's answer
2021-06-15T10:00:21-0400
"W=kq(\dfrac{1}{r_2}-\dfrac{1}{r_1}),""W=9\cdot10^9\ \dfrac{N\cdot m^2}{C^2}\cdot0.02\cdot10^{-6}\ C\cdot(\dfrac{1}{0.05\ m}-\dfrac{1}{0.4\ m}),""W=3150\ J."